OpenStudy (anonymous):

each day the wildlife commission releases 100 fish into the Durkee River in the following proportions: 55 brown trout and 45 rainbow trout. all of the fish swim through a watch station one at a time on the way upstream. what expression represents the way to calculate the probability of a passing fish being a rainbow trout?

OpenStudy (jack1):

P(rainbow) = 45/100 =9/20 =0.45 or 45% chance